OpenStudy (anonymous):

real cube roots of -64

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Second one is right this time :D

OpenStudy (anonymous):

\[\sqrt[3]{-64}= \sqrt[3]{-1}\times \sqrt[3]{64}\]\[\sqrt[3]{-1}=i\]\[\sqrt[3]{64}= \sqrt[3]{2\times2\times2\times2\times2\times2\times2}= 4\]\[i \times4=4i\]

OpenStudy (anonymous):

If I deserve medal then give me one